# SC2 Workshop 2024: Satisfiability Checking and Symbolic Computation

### v1

@@ 3.2.1 Lazy Verification
An alternative to the approach outlined above, the template of the solution does not have to be verified all at once. Instead, we could use QE to find a class of solutions for a fixed template and then prove there are no other solutions. The advantage of this approach is illustrated by the following example.
Problem C1 asks for all functions $f$ such that
@@
In this case, the solvers listed in Section 3.1 are unable to prove all solutions are a monomial quadratic (or that they must be quadratic). However, we can still apply quantifier elimination to find all monomial quadratic solutions. For this example, z3 timed out even when given 10 minutes. On the other hand, Tarski returned the formula $-1+a=0$ (yielding the solution $a=1$, i.e., $x^{2}$) in less than a second. Since we were unable to prove all solutions must belong to the monomial quadratic template class, it is still possible there are more solutions outside the monomial quadratic template class. However, now that we have the specific solution $f(x)=x^{2}$ we can ask an SMT2 solver to prove this is the only solution by assuming the equation and also the negation of $\forall x:\mathbb{R}.f(x)=x^{2}$. The solver cvc5 is easily able to show this is unsatisfiable, so that we know $f(x)=x^{2}$ is the only solution.
@@ 4 Experiments
We report results of the template-and-qe method (Section 3) on the benchmark presented in Section 2. The lazy extension of our approach (Section 3.2.1) was left for future work.
The template verification (Section 3.1) has proven to be the most difficult task. Table LABEL:tab:gencov summarizes the obtained results for template verification. For all these we were able to solve the QE task. Thirteen instances were solved completely by the automated method. Out of these thirteen, two can be traced to a competition event. Problem U24 comes from Baltic Way competition—see Section 2 for “human” solution. Problem U71 comes from The Prague seminar (PraSe-18-6-1). In both cases, the only solution is a function that is constantly 0.
The verification tasks results often in satisfiable problems, namely if there exists a solution outside of the proposed template (the $\times$ symbol in Table LABEL:tab:gencov). These satisfiable problems are nontrivial because they involve creating a counterexample to the template. For this purpose, we also ran the SyGuS approach666option --sygus-inference of cvc5 and its linear model builder [17].
We also report on the verification of the correctness of the handwritten solutions, which comprises two components: prove — show that all possible solutions are covered by the suggested solutions; check — check that all suggested solutions are indeed solutions to the problem. Table 0(b) summarizes the obtained results. Sixty-five of the provided handwritten solutions were successfully checked, i.e., the individual solutions satisfy the original specifications. Twenty of the handwritten solutions were shown to cover all the individual solutions. Unsurprisingly, the proving task is harder than the checking task. We remark that checking solutions for U79 is trivial because there are no individual solutions. More detailed results can be found on the authors’ web page. [^7] https://sat.inesc-id.pt/~mikolas/sc2_2024/
@@ 5 Conclusions and Future Work
This paper joins the effort of rising up to the challenge of making computers as powerful as are the golden medalist at the International Math Olympics (IMO). Efforts such as AlphaGeometry [18], show that machine learning models are useful for the task but at the same time, further research shows that they benefit from existing symbolic methods [19]. Here we show that symbolic methods are indeed already powerful in solving a highly nontrivial task of finding all functions fulfilling a certain specification. Besides the task leading to undecidable questions, its difficulty also lies in the fact it is not a decision problem but the response is a description of a class of mathematical objects.

### v2

@@ 3.3.1 Lazy Verification
We now briefly consider an alternative to the approach outlined above. When considering a template, we do not need to verify that all solutions are in the template class in advance. Instead, we could use QE to find a class of solutions for a fixed template and then prove there are no other solutions. The advantage of this approach is illustrated by the following example. Problem C1 asks for all functions $f$ such that the following holds
@@
In this case, the solvers listed in Section 3.1 are unable to prove that all solutions are a monomial quadratic (or that they must be quadratic). However, we can still apply quantifier elimination to find all monomial quadratic solutions. For this example, Z3 timed out even when given 10 minutes. On the other hand, Tarski returned the formula $-1+a=0$ (yielding the solution $a=1$, i.e., $x^{2}$) in less than a second. Since we were unable to prove all solutions must belong to the monomial quadratic template class, it is still possible there are more solutions outside the monomial quadratic template class. However, now that we have the specific solution $f(x)=x^{2}$ we can ask an SMT solver to prove this is the only solution by assuming the equation and also the negation of $\forall x:\mathbb{R}.f(x)=x^{2}$. The solver cvc5 is easily able to show this is unsatisfiable, so we know that $f(x)=x^{2}$ is the only solution.
@@ 4 Experiments
We report the results of the template-and-qe method (Section 3) on the benchmark presented in Section 2. The lazy extension of our approach (Section 3.3.1) was left for future work. The template verification (Section 3.1) has proven to be the most difficult task. Table LABEL:tab:gencov summarizes the obtained results for template verification. For all these, we were able to solve the QE task. Thirteen instances were solved completely by the automated method. Out of these thirteen, two can be traced to a competition event. Problem U24 comes from Baltic Way competition—see Section 2 for “human” solution. Problem U71 comes from The Prague Seminar (PraSe-18-6-1). In both cases, the only solution is a function that is constantly 0.
The verification tasks results often in satisfiable problems, namely if there exists a solution outside of the proposed template (the $\times$ symbol in Table LABEL:tab:gencov). These satisfiable problems are nontrivial because they involve creating a counterexample to the template. For this purpose, we also ran the SyGuS approach of cvc5 (option --sygus-inference) and its linear model builder [27].
We also report on the verification of the correctness of the handwritten solutions, which comprises two components: prove — show that all possible solutions are covered by the suggested solutions; check — check that all suggested solutions are indeed solutions to the problem (see Section 2). Table 0(b) summarizes the obtained results. Sixty-five of the provided handwritten solutions were successfully checked, i.e., the individual solutions satisfy the original specifications. Twenty of the handwritten solutions were shown to cover all the individual solutions. Unsurprisingly, the proving task is harder than the checking task. We remark that checking solutions for U79 is trivial because there are no individual solutions. More detailed results can be found on the authors’ web page [28].
@@ 5 Conclusions and Future Work
This paper joins the effort of rising to the challenge of making computers as powerful as the golden medalists at the International Math Olympics (IMO). Efforts such as AlphaGeometry [29], show that machine learning models are useful for the task but at the same time, further research shows that they benefit from existing symbolic methods [30]. Here we show that symbolic methods are indeed already powerful in solving a highly nontrivial task of finding all functions fulfilling a certain specification. Besides the task leading to undecidable questions, its difficulty also lies in the fact it is not a decision problem but the response is a description of a class of mathematical objects.